While passing a turn, the driver of the car applied on the steering wheel, a pair of forces with torque M = 15 Nm. What is the strength of this pair when the diameter of the steering wheel is d = 30 cm?


1
Expert's answer
2021-03-08T08:31:14-0500

The diameter of the wheel is 30 cm and the pivot point is 15 cm away from the outer ring. Both forces act in the same direction of rotation, so, the total torque is


"\tau=\frac d2F+\frac d2F=Fd,\\\space\\
F=\frac \tau d=50\text{ N}."